THE AMOUNT OF A 12% SALINE SOLUTION A LAB ASSISSTANT SHOULD ADD TO 80CC OF A 22% SALINE SOLUTION IN ORDER TO HAVE A 16% SOLUTION
Answer by mananth(16946)

--------percent ---------------- quantity
Saline I 22 ---------------- 80 cc
Saline II 12 ---------------- x cc
Mixture 16 ---------------- 80 + x cc

22 * 80 + 12 x = 16 ( 80 + x )
1760 + 12 x = 1280 + 16 x
12 x + -16 x = 1280 - -1760
-4 x = -480
/ -4
x = 120
120 cc of Saline II